OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

id like step by step help i have been having trouble for quiet a while...

OpenStudy (anonymous):

y+1=6(x-5)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i need to know how to get the answer

OpenStudy (anonymous):

ummm JAS

OpenStudy (anonymous):

hu?

OpenStudy (anonymous):

just a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m= y1-y2 divided by x1-x2

OpenStudy (anonymous):

y-y1=m(x-x1) y-(-1)=6(x-5) y+1=6(x-5) y+1=6x-30 6x-y-30-1=0 6x-y-31=0 complete ans.

OpenStudy (anonymous):

x1 is the 5 in (5,-1) and the -1 is the y1 and the slope or m is 6?

OpenStudy (anonymous):

any problem emmah

OpenStudy (anonymous):

yepper

OpenStudy (anonymous):

what is

OpenStudy (anonymous):

ok thanks i have another question im gonna post it here too

OpenStudy (anonymous):

Write the equation for the vertical line passing through (5, 8).

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

u know eq is parallel to y axis?

OpenStudy (anonymous):

im sorry my internet is bad what were you saying?

OpenStudy (anonymous):

my qestion is eqution is parallel to y axis is.

OpenStudy (anonymous):

um thats my problem im not undertsanding whst you are sayinmg....

OpenStudy (anonymous):

wht do u want and repeat the question? plzzzzz

OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.